# Evaluating the Effect of Retrieval Augmentation on Social Biases

### v1

@@ Appendix A Evaluation Metrics
When presented with ambiguous contexts where the ground-truth answer is always UNKNOWN, we calculate accuracy given by (1).
@@
Here, $n_{a}$ denotes the total number of ambiguous questions, and $n_{au}$ counts how often the model correctly responds with UNKNOWN.
@@
[…]
@@
[…]
=$=\begin{cases}2Acc_{d},&0\leq Acc_{d}\leq 0.5\\ 2(1-Acc_{d}),&0.5<Acc_{d}\leq 1\end{cases}$ (7)
Appendix B Experimental Settings
We use the following open-source LLMs in our experiments as the generator LLMs, which are available from HuggingFace: Llama3-8B, Llama3-8B-Instruct, Mistral, Mistral-Instruct, Qwen2.5-7B, Qwen2.5-7B-Instruct, Qwen2.5-3B, Qwen2.5-3B-Instruct, Qwen2.5-14B, Qwen2.5-14B-Instruct, LLM-JP-3-1.8B-Instruct, LLM-JP-3-3.7B-Instruct, LLM-JP-3-13B-Instruct. For GPT-3.5-turbo, we use the OpenAI API.
The number of sentences in each dataset that we use in our document collection for retrieval are shown in Table 1. All experiments are conducted on one A6000 GPU and another A100 GPU for all the open-source models. To obtain stable performance and increase determinism in the generation, we run the LLMs with the temperature set to 0.
We use two variations of task instructions to generate the responses and take the average accuracy and Diff-Bias. The task instructions are shown in Table 6 and the full prompt in Figure 5.
@@ Appendix B Experimental Settings
Appendix C Additional Accuracy Evaluations
@@
In this section, we report the accuracy scores for all of the experimental results that were shown in the main body of the paper using Diff-Bias scores. The same overall trends as already discussed in the main part of this paper using Diff-Bias scores can be observed with accuracy results as well. Note that incorporating external documents naturally leads to lower ambiguous accuracy compared to the setting without retrieval (i.e. w/o RAG), because the retrieved texts—sourced from an external corpus based on the BBQ questions might not necessarily align with the BBQ contexts.
§ C.1 Accuracy Across Bias Categories
@@
Table 7 reports the ambiguous and disambiguated accuracy scores for four bias categories (i.e. Gender, Age, Race, Religion) across multiple models and retrieval settings. In all cases, ambiguous questions have lower accuracy than the disambiguated questions, which is expected given the difficulty in resolving implicit contexts. Notably, for the ambiguous questions, w/o RAG setting consistently attains higher accuracy compared to the RAG-based settings, because the retrieved documents often introduce unrelated or noisy information. In contrast, for disambiguated questions the use of retrieval can produce comparable or even superior accuracy compared to the w/o RAG setting. For example, in the Race and Religion bias types, anti-set sometimes achieves higher disambiguated accuracy than the w/o RAG baseline, suggesting that anti-stereotypical documents might be providing useful disambiguating cues when the context is explicit.
§ C.2 Accuracy on the English BBQ Gender Dataset
Table 8 shows the accuracy scores on the English BBQ dataset across different corpus settings and a range of models. Consistent with the observations above, ambiguous questions generally exhibit the highest accuracy in the w/o RAG setting. For instance, GPT-3.5 achieves an accuracy of 45.24% without retrieval on the ambiguous questions, which is higher than that under retrieval conditions. Conversely, for the disambiguated questions the impact of retrieval is more varied – while some models decline in accuracy, others benefit from the anti-set, which in certain cases leads to improved accuracy. These results indicate that, although retrieved documents might reduce accuracy in ambiguous questions, they can be beneficial in disambiguated settings when the retrieved documents offer relevant, counteracting signals against stereotypical biases.
§ C.3 Multi-lingual Accuracy Evaluations
Table 9 presents the accuracy for Chinese (CBBQ) and Japanese (JBBQ) datasets. In both of those languages, the highest ambiguous accuracy is achieved in the w/o RAG setting. When RAG is applied, the full-set reports the highest ambiguous accuracy, while the anti-set generally results in the lowest ambiguous accuracy. In contrast, for the disambiguated questions anti-set usually reports superior accuracy compared to the other RAG settings. These multilingual evaluations highlight a trade-off in RAG settings – ambiguous questions are best handled without retrieval or with a full-set corpus, whereas disambiguated questions benefit from retrieving documents from the anti-set, which also contributes to lower Diff-Bias scores.
§ C.4 Effect of the Retrievers on Accuracy
@@
§ C.5 Effect of Varying the Number of Retrieved Documents on Accuracy
@@
Figure 6 and Figure 7 compare the accuracy of three LLMs under different numbers of retrieved documents. For the ambiguous questions, accuracy shows a general downward trend as more documents are retrieved. Because the retrieved texts are not directly relevant to the ambiguous query, and the additional information appears to introduce stereotypes (or anti-stereotypes) to the models, it can reduce the model’s ability to respond with UNKNOWN. By contrast, for the disambiguated questions, retrieving more documents sometimes achieve accuracy that is comparable (or at times exceeds) to the w/o RAG setting.
Bias Category Setting GPT-3.5 Llama3-8B-Inst. Qwen2.5-7B-Inst. Qwen2.5-14B Qwen2.5-14B-Inst.
Gender w/o RAG 45.24 / 75.74 50.03 / 60.52 81.45 / 52.03 63.79 / 82.84 96.53 / 71.92
stereo-set 27.83 / 71.68 26.19 / 63.74 62.6 / 53.72 46.83 / 77.63 87.05 / 68.30
full-set 27.58 / 73.12 24.36 / 63.89 62.95 / 56.10 49.75 / 77.08 89.78 / 67.76
anti-set 22.07 / 72.97 23.66 / 66.07 58.09 / 58.09 41.02 / 78.57 83.63 / 68.90
Age w/o RAG 18.97 / 88.7 31.03 / 75.57 60.08 / 77.42 42.80 / 92.53 78.76 / 89.24
stereo-set 16.63 / 81.55 16.03 / 70.03 48.64 / 77.99 35.30 / 89.65 75.95 / 83.32
full-set 19.97 / 83.21 15.76 / 71.47 51.17 / 79.59 40.84 / 90.43 87.93 / 84.51
anti-set 16.44 / 84.35 14.57 / 71.49 50.11 / 78.61 37.42 / 90.46 87.36 / 84.89
Race w/o RAG 56.97 / 83.62 59.04 / 77.55 94.04 / 68.03 80.85 / 93.62 98.94 / 78.46
stereo-set 33.88 / 83.24 35.00 / 78.03 70.32 / 75.85 58.35 / 92.66 91.33 / 83.83
full-set 35.74 / 85.16 37.61 / 78.14 73.40 / 76.06 62.71 / 94.04 94.04 / 84.15
anti-set 35.21 / 86.44 38.94 / 80.69 79.95 / 74.04 55.64 / 94.95 96.81 / 86.65
Religion w/o RAG 49.08 / 80.17 60.67 / 74.25 84.58 / 64.25 67.75 / 83.67 90.33 / 69.42
stereo-set 37.92 / 77.08 38.67 / 75.42 71.67 / 68.92 53.05 / 87.67 87.25 / 68.5
full-set 35.67 / 78.67 35.50 / 74.25 66.5 / 70.67 51.50 / 86.25 86.75 / 69.92
anti-set 30.50 / 78.58 32.92 / 76.25 67.17 / 71.75 47.67 / 87.92 84.42 / 72.17
@@
Table 7: Accuracy scores for the ambiguous and disambiguated contexts (separated by ‘/’) for different bias categories and models, when document collections with varying degrees of social biases are used for retrieval. In each sub-category (Gender, Age, Race, Religion), the scores for each model are compared vertically. For each model and bias category, the maximum value in the ambiguous (left) and disambiguated (right) scores is highlighted in light red bold, while the minimum is highlighted in light blue bold.
@@
GPT-3.5 45.24 / 75.74 27.83 / 71.68 27.58 / 73.12 22.07 / 72.97
Llama3-8B 25.94 / 41.96 21.03 / 47.97 21.43 / 47.82 20.78 / 49.40
Llama3-8B-Inst. 50.30 / 60.52 26.19 / 63.74 24.36 / 63.89 23.66 / 66.07
Mistral 16.12 / 54.02 19.69 / 50.84 17.71 / 49.45 18.40 / 49.85
Mistral-Inst. 66.91 / 67.26 45.49 / 69.25 47.22 / 71.38 42.11 / 71.88
Llm-jp-1.8B 7.04 / 48.21 16.96 / 43.75 18.06 / 44.25 16.27 / 45.85
Llm-jp-3.7B 10.52 / 51.49 18.65 / 47.52 19.35 / 46.23 16.27 / 45.14
Llm-jp-13B 10.62 / 82.74 6.75 / 78.27 7.14 / 78.27 6.75 / 77.78
Qwen-7B 30.65 / 67.31 26.49 / 68.40 26.74 / 69.54 24.31 / 69.05
Qwen-7B-Inst. 81.45 / 52.03 62.60 / 53.72 62.95 / 56.10 58.09 / 58.09
Qwen-3B 8.23 / 78.97 4.12 / 76.49 3.22 / 76.24 2.88 / 77.83
Qwen-3B-Inst. 68.20 / 58.43 57.19 / 57.14 52.28 / 63.10 54.37 / 59.42
Qwen-14B 63.79 / 82.84 46.83 / 77.63 49.75 / 77.08 41.02 / 78.57
Qwen-14B-Inst. 96.53 / 71.92 87.05 / 68.30 89.78 / 67.76 83.63 / 68.90
Table 8: Comparison of accuracy scores across different corpus settings on the BBQ gender dataset. Scores are reported in the format ambiguous / disambiguous, where higher values indicate better performance. For each model, the maximum ambiguous and disambiguous scores are highlighted in light red bold, while the minimum values are highlighted in light blue bold.
Model CBBQ JBBQ
w/o RAG stereo-set full-set anti-set w/o RAG stereo-set full-set anti-set
GPT-3.5 26.52 / 64.30 13.31 / 67.08 16.77 / 67.28 12.45 / 68.42 30.52 / 52.68 23.31 / 55.93 27.40 / 56.08 24.26 / 56.29
Qwen-7B-Inst. 90.48 / 45.88 37.55 / 64.09 43.40 / 62.04 35.50 / 64.30 77.56 / 53.66 48.29 / 57.54 50.08 / 58.21 45.35 / 58.00
Qwen-14B 72.62 / 57.30 42.42/ 60.29 49.46 / 61.11 44.05 / 61.52 42.56 / 77.89 28.71 / 73.90 31.06 / 75.23 25.95 / 77.28
Qwen-14B-Inst. 96.32 / 40.84 76.73 / 45.78 84.52 / 48.97 75.97 / 46.09 82.31 / 78.35 61.84 / 77.86 67.15 / 78.20 61.61 /80.52
Table 9: Accuracy scores for Chinese (CBBQ) and Japanese (JBBQ) datasets. Accuracy values are reported in the format ambiguous / disambiguated, where higher values indicate better performance. For each model, the maximum ambiguous and disambiguated scores are highlighted in light red bold, while the minimum values are highlighted in light blue bold.

### v2

@@ Appendix A Evaluation Metrics
When presented with ambiguous contexts where the ground-truth answer is always “UNKNOWN”, we calculate accuracy given by (1).
@@
Here, $n_{a}$ denotes the total number of ambiguous questions, and $n_{au}$ counts how often the model correctly responds with “UNKNOWN”.
@@
For the ambiguous contexts, the Diff-Bias score, $\text{Diff-bias}_{a}$, is defined as the difference between the proportion of the stereotypical answers and counter-stereotypical answers, as given by (3).
$\text{Diff-bias}_{a}=\frac{n_{as}-n_{ac}}{n_{a}}$ (3)
@@
$|\text{Diff-bias}_{a}|\leq 1-Acc_{a},\quad(0\leq Acc_{a}\leq 1)$ (4)
An unbiased model would have $\text{Diff-bias}_{a}=0$, while a model that consistently favours stereotypical responses would return $\text{Diff-bias}_{a}=1$ (or 100 when expressed as a percentage).
For the disambiguated contexts, the Diff-Bias score, $\text{Diff-bias}_{d}$, is defined as the difference between the accuracy on the stereotyped contexts (${\rm Acc}_{ds}$) and the accuracy on counter-stereotyped contexts (${\rm Acc}_{dc}$) as given by (5).
$\text{Diff-bias}_{d}=Acc_{ds}-Acc_{dc}=\frac{n_{ss}}{n_{s}}-\frac{n_{cc}}{n_{c}}$ (5)
Here, $n_{ss}$ and $n_{cc}$ are the correctly answered instances in stereotyped and counter-stereotyped contexts, respectively, and $n_{s}$ and $n_{c}$ represent the total number of each type of contexts. The range of $\text{Diff-bias}_{d}$ is given by (6).
$|\text{Diff-bias}_{d}|\leq 1-|2Acc_{d}-1|,\quad(0\leq Acc_{d}\leq 1)$ (6)
$=\begin{cases}2Acc_{d},&0\leq Acc_{d}\leq 0.5\\ 2(1-Acc_{d}),&0.5<Acc_{d}\leq 1\end{cases}$ (7)
Appendix B Translated Documents Human Examination
To assess the extent to which machine translation preserves social biases, we randomly sampled 100 documents per target language (50 stereotypical, 50 anti-stereotypical) and presented only the translated versions for evaluation. Each translation was independently scored by two native speakers of the target language in our department who share the same cultural background described in the source texts.
We employed a four-point scale to rate semantic agreement between source sentences and their target translations (higher the better): 1 = unrelated (the meaning in the source sentence is lost in the translation), 2 = weakly related (the main idea from the source sentence is present in the translation to an extent but not fully), 3 = strongly related (most content in the source sentence is preserved in the translation, while minor issues such as incorrect/missing gender pronouns can be observed), 4 = fully faithful (original meaning of the source sentence is completely conveyed by the translated sentence). Table 8 reports the average distribution of human scores for each language and stereotype type.
Overall, Chinese translations achieved slightly higher translation agreement than Japanese. Annotator feedback revealed that 16% of the Japanese translations omitted gender pronouns – compared to only 2% for Chinese – which hinder the retrieval of relevant documents in the JBBQ dataset.
Rating 4 3 2 1
Chinese 90 6 4 0
Japanese 79 17.5 2.5 1
Table 8: Human rating percentage distribution for Chinese and Japanese translations when source sentences are given in English.
Appendix C Experimental Settings
We use the following open-source LLMs in our experiments as the generator LLMs, which are available from HuggingFace: Llama3-8B, Llama3-8B-Instruct, Mistral, Mistral-Instruct, Qwen2.5-7B, Qwen2.5-7B-Instruct, Qwen2.5-3B, Qwen2.5-3B-Instruct, Qwen2.5-14B, Qwen2.5-14B-Instruct, Qwen2.5-72B-Instruct, LLM-JP-3-1.8B-Instruct, LLM-JP-3-3.7B-Instruct, LLM-JP-3-13B-Instruct. For GPT-3.5-turbo and O4-mini, we use the OpenAI API.
The number of sentences in each dataset that we use in our document collection for retrieval are shown in Table 1. To ensure that the generator would not be affected by being trained on a dataset with a limited set of specific proper names, we follow the Huang and Xiong (2024) and remove all proper names from the benchmarks.
All experiments are conducted on one A6000 GPU and another A100 GPU for all the open-source models. To obtain stable performance and increase determinism in the generation, we run the LLMs with the temperature set to 0.
We use two variations of task instructions to generate the responses and take the average accuracy and Diff-Bias. The task instructions are shown in Table 9 and the full evaluation prompt is shown in Figure 5.
@@ Appendix C Experimental Settings
Appendix D Intersectional Biases
Prior work studying social biases in LLMs has shown that social bias categories are not necessarily independent and there often exist correlations between different bias categories (Ma et al., 2023; Tan and Celis, 2019; Riedel and Rau, 2025; Lalor et al., 2022). Tan and Celis (2019) studied the intersectional bias between gender and race categories, and reported such intersectional biases to exist in BERT and GPT-2 models. Riedel and Rau (2025) showed that race and religion are highly correlated and their interplay influences political decisions. Lalor et al. (2022) conducted an extensive analysis covering multiple models and debiasing methods, and showed intersectional biases across multiple social bias types. Ma et al. (2023) proposed a benchmark dataset for evaluating intersectional social biases. However, to the best of our knowledge, no prior work has studied how intersectional biases are influenced under RAG, which we aim to study in this section.
Specifically, we study the intersectional biases when a query to the RAG system expresses a social bias type, which does not appear in the indexed document collection. Recall that our document collection (statistics provided in Table 1) consists of four social bias types: gender, age, race and religion. To evaluate how intersectional biases are affected under RAG, we select ambiguous and disambiguated questions from the English BBQ dataset for four additional social bias categories: Sexual Orientation, Physical Appearance, Nationality and Disability. We then compare social biases in GPT-3.5-turbo (closed source) and Qwen2.5-7B-Inst. (open source) using Diff-Bias scores of the original models (i.e. w/o RAG) vs. when the stereo-set was used as the document collection for RAG.
Table 10 shows the resulting Diff‐Bias scores for the ambiguous and disambiguated contexts. We see that compared to the w/o RAG baseline, for every bias type, retrieving documents outside the target category still increases the Diff‐Bias scores, except for Disability for GPT-3.5 in the disambiguated contexts. This shows that intersectional social biases are amplified under RAG. Moreover, we see a high degree of bias categories between queries and the retrieved documents. For example, among the top-10 documents retrieved for Nationality questions, 67.95% were drawn from the Race category; for Sexual Orientation questions, 88.63% came from Gender‐stereotypical documents—yet both cases show an amplified bias score even though the retrieved texts did not match the query’s own bias type. This is an interesting and novel finding that further emphasizes the seriousness of the bias amplification issue under RAG because a stereotypical document collection can amplify not only the social biases contained in the collection but also ones that are not.
Bias Type Setting GPT-3.5-turbo Qwen-7B-Inst.
Sexual Orientation w/o RAG 6.71 / -2.31 5.21 / -6.71
Stereo Set 7.06 / -1.62 5.21 / -1.15
Physical Appearance w/o RAG 26.33 / 12.94 10.79 / -0.25
Stereo Set 30.07 / 14.21 23.54 / 2.92
Nationality w/o RAG 17.75 / 2.08 11.82 / 3.18
Stereo Set 24.09 / 10.71 16.98 / 2.86
Disability w/o RAG 23.14 / 10.53 3.41/-0.25
Stereo Set 23.39 / 9.51 7.52/2.44
Table 10: Diff-Bias scores for the ambiguous and disambiguated contexts (separated by ‘/’) for different bias types which are not covered by the retrieval documents.
Figure 6: The prompt use to evaluate the Bias-Level Scores in the retrieved sentence.
Appendix E Relationship between number of retrieved documents and their Social Biases
To further study the trend observed in § 4.4 where we hypothesised that there might be a trade-off between the relevance of the retrieved documents and the effect on the Diff-Bias scores, we conducted a detailed analysis as follows. Specifically, we use the anti-set with disambiguated questions, varying the number of retrieved documents ($k$). We define two key metrics for each retrieved set: a relevance score and a bias score as described next.
  • Relevance Score. For a given query and a set of $k$ retrieved documents, the relevance score was calculated as the average cosine similarity between the query embedding and the embedding of each document in the set. The higher the semantic similarity of a document to the query, it is assumed that the document to be more relevant to the query.
  • Bias-Level Score. To measure the level of bias in the retrieved documents, we adopted a method proposed by (Kumar et al., 2024), who demonstrated a high correlation between LLM-based scores and human evaluations of gender bias. We employed GPT-4o as an annotator to score each retrieved document on a scale from -3 (strongly anti-stereotype) to +3 (strongly stereotype). The prompt used for this purpose is shown in Figure 6. The average of these individual scores was then taken as the final bias-level score for the entire retrieved set.
To evaluate the agreement between LLM-based bias-level ratings and that by humans, we conducted a pilot study where we randomly sampled 60 cases and asked three linguistically trained human annotators from our department to annotate the bias-level using the same instruction as provided to the LLM. We then computed the average Spearman’s Rank Correlation coefficient between the LLM-provided ratings and the human annotated ratings. We obtained an average Spearman’s correlation coefficient of 0.690, which indicates a moderate positive correlation between human annotations and the LLM ratings.
$k$ Relevance Bias-Level Harmonic Mean
3 0.846 -0.198 0.320
5 0.842 -0.266 0.405
10 0.835 -0.384 0.526
20 0.829 -0.377 0.518
30 0.818 -0.384 0.523
Table 11: Analysis of Relevance vs. Bias-Level for the anti-set.
As shown in Table 11, the relevance scores show a monotonic decrease as $k$ increases, confirming that top-ranked documents are indeed the most semantically relevant to the query. In contrast, the bias-level scores first increase and then drop as we go down the ranked list of documents. Specifically, the magnitude of the anti-stereotype signal (the absolute value of the score) increases from $k=3$, peaks at $k=10$, after which it begins to slightly drop. This suggests that the top ranked documents may not be the strongest anti-stereotypical documents.
To quantitatively model this trade-off, we calculated the harmonic mean of the relevance score and the absolute bias-level score. The result, which can be interpreted as a combined Relevance-Bias Impact Score, peaks at k=10 (0.526), perfectly aligning with the trend observed in the Diff-Bias scores (see Figure 4, right).
Appendix F Additional Accuracy Evaluations
@@
In this section, we report the accuracy scores for all of the experimental results that were shown in the main body of the paper using Diff-Bias scores. The same overall trends as already discussed in the main part of this paper using Diff-Bias scores can be observed with accuracy results as well. Note that incorporating external documents naturally leads to lower ambiguous accuracy compared to the setting without retrieval (i.e. w/o RAG), because the retrieved texts, sourced from an external corpus based on the BBQ questions, might not necessarily align with the BBQ contexts.
§ F.1 Accuracy Across Bias Categories
@@
Table 12 reports the ambiguous and disambiguated accuracy scores for the four bias categories (i.e. Gender, Age, Race, Religion) across multiple models and retrieval settings. In all cases, ambiguous questions have a lower accuracy than the disambiguated questions, which is expected given the difficulty in resolving implicit contexts. Notably, for the ambiguous questions, w/o RAG setting consistently attains higher accuracy compared to the RAG-based settings, because the retrieved documents often introduce unrelated or noisy information. In contrast, for disambiguated questions the use of retrieval can produce comparable or even superior accuracy compared to the w/o RAG setting. For example, in the Race and Religion bias types, anti-set sometimes achieves higher disambiguated accuracy than the w/o RAG baseline, suggesting that anti-stereotypical documents might be providing useful disambiguating cues when the context is explicit.
§ F.2 Accuracy on the English BBQ Gender Dataset
Table 13 shows the accuracy scores on the English BBQ dataset across different corpus settings and a range of models.
A consistent trend is that retrieval, regardless of the document polarity, degrades accuracy on ambiguous questions. This suggests that the presence of external context—even neutral or anti-stereotypical—can mislead models into making a definitive choice when “UNKNOWN” is the correct answer. For example, the accuracy of Mistral-Inst. drops significantly from 66.91% in the w/o RAG setting to 45.49% with the stereo-set. This indicates a heightened risk of RAG systems producing seemingly confident but baseless answers in information-poor scenarios.
Conversely, for disambiguated questions, the impact of retrieval is more nuanced. While stereotypical documents often cause a minor decline in accuracy, the anti-set can be beneficial. Several models, including Llama3-8B-Inst. (60.52% to 66.07%) and Mistral-Inst. (67.26% to 71.88%), show improved accuracy with anti-stereotypical documents.
Regarding model scale, larger models generally exhibit higher accuracy. The accuracy of even the largest models, such as Qwen-72B-Inst., drops substantially on ambiguous questions when using the stereo-set (from 98.26% to 86.21%). The reasoning model, o4-mini, is notably more robust. On disambiguated questions, its accuracy remains exceptionally high and stable (96%) across all settings, indicating that its strong reasoning capabilities allow it to reliably prioritise the explicit evidence in the prompt over any distracting retrieved context.
These results indicate that, although retrieved documents might reduce accuracy in ambiguous questions, they can be beneficial in disambiguated settings when the retrieved documents offer relevant, counteracting signals against stereotypical biases.
§ F.3 Effect of the Retrievers on Accuracy
@@
§ F.4 Effect of Varying the Number of Retrieved Documents on Accuracy
@@
Figure 7 and Figure 8 compare the accuracy of three LLMs under different numbers of retrieved documents. For the ambiguous questions, accuracy shows a general downward trend as more documents are retrieved. Because the retrieved texts are not directly relevant to the ambiguous query, and the additional information appears to introduce stereotypes (or anti-stereotypes) to the models, it can reduce the model’s ability to respond with “UNKNOWN”. By contrast, for the disambiguated questions, retrieving more documents sometimes achieve accuracy that is comparable (or at times exceeds) to the w/o RAG setting.
§ F.5 Multi-lingual Accuracy Evaluations
Table 15 and Table 16 present the accuracy for Chinese (CBBQ) and Japanese (JBBQ) datasets. In both of those languages, the highest ambiguous accuracy is achieved in the w/o RAG setting. When RAG is applied, the anti-set generally results in the lowest ambiguous accuracy. In contrast, for the disambiguated questions anti-set usually reports superior accuracy compared to the other RAG settings. These multilingual evaluations suggests a potential trade-off in RAG settings – ambiguous questions are best handled without retrieval or with a full-set corpus, whereas disambiguated questions benefit from retrieving documents from the anti-set, which also contributes to lower Diff-Bias scores.
Bias Category Setting GPT-3.5 Llama3-8B-Inst. Qwen2.5-7B-Inst. Qwen2.5-14B Qwen2.5-14B-Inst. Qwen2.5-72B-Inst.
Gender w/o RAG 45.24±2.17 / 75.74±1.87 50.03±2.18 / 60.52±2.13 81.45±1.70 / 52.03±2.18 63.79±2.10 / 82.84±1.65 96.53±0.80 / 71.92±1.96 98.26±0.51 / 68.25±2.03
stereo-set 27.83±1.96 / 71.68±1.97 26.19±1.92 / 63.74±2.10 62.60±2.11 / 53.72±2.18 46.83±2.18 / 77.63±1.82 87.05±1.47 / 68.30±2.03 86.21±1.38 / 68.61±2.02
full-set 27.58±1.95 / 73.12±1.94 24.36±1.87 / 63.89±2.10 62.95±2.11 / 56.10±2.17 49.75±2.18 / 77.08±1.83 89.78±1.32 / 67.76±2.04 89.44±1.23 / 65.97±2.07
anti-set 22.07±1.81 / 72.97±1.94 23.66±1.86 / 66.07±2.07 58.09±2.15 / 58.09±2.15 41.02±2.15 / 78.57±1.79 83.63±1.62 / 68.90±2.02 84.13±1.50 / 69.40±2.01
Age w/o RAG 18.97±1.27 / 88.70±1.02 31.03±1.49 / 75.57±1.39 60.08±1.58 / 77.42±1.35 42.80±1.60 / 92.53±0.85 78.76±1.32 / 89.24±1.00 84.41±1.48 / 91.93±1.14
stereo-set 16.63±1.20 / 81.55±1.25 16.03±1.19 / 70.03±1.48 48.64±1.61 / 77.99±1.34 35.30±1.54 / 89.65±0.98 75.95±1.38 / 83.32±1.20 73.18±1.78 / 91.50±1.20
full-set 19.97±1.29 / 83.21±1.21 15.76±1.18 / 71.47±1.46 51.17±1.62 / 79.59±1.30 40.84±1.59 / 90.43±0.95 87.93±1.05 / 84.51±1.17 84.62±1.48 / 90.08±1.29
anti-set 16.44±1.20 / 84.35±1.17 14.57±1.14 / 71.49±1.46 50.11±1.62 / 78.61±1.32 37.42±1.56 / 90.46±0.95 87.36±1.07 / 84.89±1.16 80.44±1.58 / 91.20±1.22
Race w/o RAG 56.97±2.24 / 83.62±1.67 59.04±2.22 / 77.55±1.89 94.04±1.07 / 68.03±2.11 80.85±1.78 / 93.62±1.10 98.94±0.46 / 78.46±1.86 99.10±0.27 / 83.83±1.58
stereo-set 33.88±2.14 / 83.24±1.69 35.00±2.16 / 78.03±1.87 70.32±2.07 / 75.85±1.93 58.35±2.23 / 92.66±1.18 91.33±1.27 / 83.83±1.66 93.78±0.81 / 82.55±1.63
full-set 35.74±2.17 / 85.16±1.61 37.61±2.19 / 78.14±1.87 73.40±2.00 / 76.06±1.93 62.71±2.19 / 94.04±1.07 94.04±1.05 / 84.15±1.65 96.23±0.69 / 80.00±1.74
anti-set 35.21±2.16 / 86.44±1.55 38.94±2.20 / 80.69±1.78 79.95±1.81 / 74.04±1.98 55.64±2.25 / 94.95±0.99 96.81±0.79 / 86.65±1.54 96.49±0.69 / 83.04±1.63
Religion w/o RAG 49.08±2.83 / 80.17±2.26 60.67±2.76 / 74.25±2.47 84.58±2.04 / 64.25±2.71 67.75±2.64 / 83.67±2.09 90.33±1.67 / 69.42±2.61 93.17±1.08 / 74.50±1.86
stereo-set 37.92±2.75 / 77.08±2.38 38.67±2.76 / 75.42±2.44 71.67±2.55 / 68.92±2.62 53.05±2.82 / 87.67±1.86 87.25±1.89 / 68.50±2.63 89.59±1.28 / 70.67±1.98
full-set 35.67±2.71 / 78.67±2.32 35.50±2.71 / 74.25±2.47 66.50±2.67 / 70.67±2.58 51.50±2.83 / 86.25±1.95 86.75±1.92 / 69.92±2.59 88.75±1.32 / 72.08±1.95
anti-set 30.50±2.61 / 78.58±2.32 32.92±2.66 / 76.25±2.41 67.17±2.66 / 71.75±2.55 47.67±2.83 / 87.92±1.84 84.42±2.05 / 72.17±2.54 87.34±1.37 / 76.09±1.86
@@
Table 12: Accuracy scores for the ambiguous and disambiguated contexts (separated by ‘/’) for different bias categories and models, when document collections with varying degrees of social biases are used for retrieval. In each sub-category (Gender, Age, Race, Religion), the scores for each model are compared vertically with 95% CI half-widths shown as subscripts.
@@
GPT-3.5 45.24±2.17 / 75.74±1.87 27.83±1.96 / 71.68±1.97 27.58±1.95 / 73.12±1.94 22.07±1.81 / 72.97±1.94
Llama3-8B 25.94±1.91 / 41.96±2.15 21.03±1.78 / 47.97±2.18 21.43±1.79 / 47.82±2.18 20.78±1.77 / 49.40±2.18
Llama3-8B-Inst. 50.30±2.18 / 60.52±2.13 26.19±1.92 / 63.74±2.10 24.36±1.87 / 63.89±2.10 23.66±1.86 / 66.07±2.07
Mistral-Inst. 66.91±2.05 / 67.26±2.05 45.49±2.17 / 69.25±2.01 47.22±2.18 / 71.38±1.97 42.11±2.16 / 71.88±1.96
Llm-jp-1.8B 7.04±1.12 / 48.21±2.18 16.96±1.64 / 43.75±2.17 18.06±1.68 / 44.25±2.17 16.27±1.61 / 45.85±2.18
Llm-jp-3.7B 10.52±1.34 / 51.49±2.18 18.65±1.70 / 47.52±2.18 19.35±1.72 / 46.23±2.18 16.27±1.61 / 45.14±2.17
Llm-jp-13B 10.62±1.34 / 82.74±1.65 6.75±1.10 / 78.27±1.80 7.14±1.12 / 78.27±1.80 6.75±1.10 / 77.78±1.81
Qwen-3B 8.23±1.20 / 78.97±1.78 4.12±0.87 / 76.49±1.85 3.22±0.77 / 76.24±1.86 2.88±0.73 / 77.83±1.81
Qwen-3B-Inst. 68.20±2.03 / 58.43±2.15 57.19±2.16 / 57.14±2.16 52.28±2.18 / 63.10±2.11 54.37±2.17 / 59.42±2.14
Qwen-7B 30.65±2.01 / 67.31±2.05 26.49±1.93 / 68.40±2.03 26.74±1.93 / 69.54±2.01 24.31±1.87 / 69.05±2.02
Qwen-7B-Inst. 81.45±1.70 / 52.03±2.18 62.60±2.11 / 53.72±2.18 62.95±2.11 / 56.10±2.17 58.09±2.15 / 58.09±2.15
Qwen-14B 63.79±2.10 / 82.84±1.65 46.83±2.18 / 77.63±1.82 49.75±2.18 / 77.08±1.83 41.02±2.15 / 78.57±1.79
Qwen-14B-Inst. 96.53±0.80 / 71.92±1.96 87.05±1.47 / 68.30±2.03 89.78±1.32 / 67.76±2.04 83.63±1.62 / 68.90±2.02
Qwen-72B 70.49±1.73 / 83.29±1.59 62.21±1.71 / 83.98±1.56 63.84±1.82 / 82.29±1.64 57.05±1.88 / 85.12±1.51
Qwen-72B-Inst. 98.26±0.51 / 68.25±2.03 86.21±1.38 / 68.61±2.02 89.44±1.23 / 65.97±2.07 84.13±1.50 / 69.40±2.01
O4-mini 95.88±0.85 / 95.53±0.89 86.45±1.30 / 96.33±0.82 86.02±1.42 / 96.48±0.79 80.31±1.60 / 96.19±0.83
Table 13: Comparison of accuracy scores across different corpus settings on the BBQ gender dataset. Scores are reported in the format ambiguous / disambiguated, where higher values indicate better performance. For each model, the 95% CI half-widths are shown as subscripts